### v1

@@ 4 Conclusion and perspectives
Datasets on blazar flares presently available with simultaneous X-ray and VHE observations with good time coverage are still very rare, and significant time delays or accurate hysteresis patterns at VHE are still difficult to obtain even for exceptional events (see e.g. Abramowski et al., 2012). However, observations at lower energies, particularly in X-rays, suggest that a majority of blazar flares tend to be in the cooling-driven regime with a decreasing trend of intrinsic delays at high fluxes (Takahashi et al., 1996; Kirk et al., 1998; Tramacere et al., 2009; Abeysekara et al., 2017; Böttcher & Baring, 2019; Abe et al., 2024), while a few cases are also observed in the acceleration-driven regime with an intrinsic delay increasing trend (Ravasio et al., 2004; Aharonian et al., 2009). Since observational constraints are very stringent on superluminal LIV effects (see e.g. Cao et al., 2021; Addazi et al., 2022), it is fair to restrict the present discussion to subluminal effects. The latter would oppose to the intrinsic trend of the cooling-driven family of flares and tend to suppress the time delays observed at VHE or send them towards the increasing trend. In such a case, analysis of simultaneous data in X-rays and gamma-rays could lead to the detection of a clockwise (respectively counter-clockwise) orientation of the hysteresis pattern in X-rays (respectively VHE gamma-rays) and the measurement of large Euclidean distances, which would provide prime clues for a LIV signal.
@@
It is important to emphasise that intrinsic delays as well as Euclidean distance and HIDs strongly constrain leptonic models, in addition to what has mostly been done with light curves and SEDs only so far. Using leptonic models including parameterised LIV effects, it could also be possible, in principle, to remove degeneracy between LIV and intrinsic delays by identifying the scenario which fits best the whole set of observable quantities. At first glance, intrinsic one-zone SSC flares, assuming that they can be firmly identified as such by multi-wavelength and multi-messenger observations, will be the most appropriate for LIV versus intrinsic delay disentangling, due to their limited number of free parameters. Indeed, the correlation between X-rays and VHE gamma-rays often observed during BL Lac flaring states has been a long standing argument in favour of SSC emission (see e.g. Sol et al., 2013), and such one-zone SSC framework still appear to provide a reasonable first order description, in the hard X-ray and VHE ranges, of some of the brightest and best observed flares detected in archetypal high-frequency peaked BL Lac objects such as PKS 2155-304, Mkn 501, and Mkn 421 (Abramowski et al., 2012; Hovatta et al., 2015; Ahnen et al., 2018; Dmytriiev et al., 2021; Abe et al., 2024). Although rare events, the development of efficient alert networks and real-time analysis, the use of data-driven detection techniques, deep learning and anomaly detection, should promote better anticipation of the onset of flares and make it possible to monitor and characterise more and more of them.
@@
However, the constantly increasing quality and complexity of the multi-wavelength datasets gathered by new generation instruments constrain the flare scenarios more and more and it will be necessary in the future to go beyond the simplistic one-zone SSC paradigm. In particular, the investigation of additional intrinsic effects in multi-zone scenarios, which would require to describe and constrain light travel effects, and in hadronic or lepto-hadronic models (Sol & Zech, 2022) will be a crucial development. The evaluation of Euclidean distance and HID methods in these more complex scenarios will still be needed. However, multi-zones scenarios and hadronic processes might face difficulties to reproduce fast-varying flares. Beyond these possibilities, alternative interesting proposals suggest that some flares could be due to pure geometrical phenomena inducing sudden changes in the Doppler factor of the emitting zone. Such models were developed in the literature for a few blazars showing periodicities or helical jet structures (Mohan & Mangalam, 2015; Raiteri et al., 2017; Sobacchi et al., 2017; Sarkar et al., 2020; Roy et al., 2022; Abe et al., 2024; Chen et al., 2024). Such type of flares would certainly deserve specific studies in the context of LIV searches, since they could provide an initial signal with possibly limited intrinsic features, or at least different from standard SSC flares.
@@
Eventually, future VHE and X-ray instruments will shed new light on this crucial quest. In particular, the development of multi-wavelength and multi-messenger astronomy and the advent of the Cherenkov Telescope Array (CTA) at VHE will provide for the first time a significant sample of blazar flares observed with good spectral and temporal resolutions and coverage, at different redshifts. This should firmly constrain the true nature of blazar flares, allow discriminating intrinsic energy-dependent time delays from LIV-induced ones, and usher in a new era for both LIV and flare modelling research.

### v2

@@ 4 Conclusion and perspectives
Datasets on blazar flares presently available with simultaneous X-ray and VHE observations with good time coverage are still very rare, and significant time delays or accurate hysteresis patterns at VHE are still difficult to obtain, even for exceptional events (see e.g. Abramowski et al., 2012). However, observations at lower energies, particularly in X-rays, suggest that the majority of blazar flares tend to be in the cooling-driven regime with a decreasing trend of intrinsic delays at high fluxes (Takahashi et al., 1996; Kirk et al., 1998; Tramacere et al., 2009; Abeysekara et al., 2017; Böttcher & Baring, 2019; Abe et al., 2024), while a few cases are also observed in the acceleration-driven regime with an intrinsic delay increasing trend (Ravasio et al., 2004; Aharonian et al., 2009). Since observational constraints are very stringent on superluminal LIV effects (see e.g. Cao et al., 2021; Addazi et al., 2022), it is fair to restrict the present discussion to subluminal effects. The latter would be opposite to the intrinsic trend of the cooling-driven family of flares and would tend to suppress the time delays observed at VHE or send them towards the increasing trend. In this case, an analysis of simultaneous data in X-rays and gamma-rays could lead to the detection of a clockwise (respectively anticlockwise) orientation of the hysteresis pattern in X-rays (respectively VHE gamma-rays) and the measurement of large Euclidean distances, which would provide prime clues for a LIV signal.
@@
It is important to emphasise that intrinsic delays as well as Euclidean distance and HIDs strongly constrain leptonic models, in addition to what has to date mostly been done only with light curves and SEDs. Using leptonic models including parametrised LIV effects, it could also be possible, in principle, to remove degeneracy between LIV and intrinsic delays by identifying the scenario which fits best the whole set of observable quantities. At first glance, intrinsic one-zone SSC flares, assuming that they can be firmly identified as such by multi-wavelength and multi-messenger observations, will be the most appropriate for LIV versus intrinsic delay disentangling, due to their limited number of free parameters. The correlation between X-rays and VHE gamma-rays often observed during BL Lac flaring states has been a long standing argument in favour of SSC emission (see e.g. Sol et al., 2013), and such a one-zone SSC framework still appears to provide a reasonable first-order description, in the hard X-ray and VHE ranges, of some of the brightest and best observed flares detected in archetypal high-frequency peaked BL Lac objects, such as PKS 2155-304, Mkn 501, and Mkn 421 (Abramowski et al., 2012; Hovatta et al., 2015; Ahnen et al., 2018; Dmytriiev et al., 2021; Abe et al., 2024). Although they are rare events, the development of efficient alert networks and real-time analysis, and the use of data-driven detection techniques (e.g. deep learning, and anomaly detection) should promote better anticipation of the onset of flares and make it possible to monitor and characterise more of them.
@@
However, the constantly increasing quality and complexity of the multi-wavelength datasets gathered by new generation instruments constrain the flare scenarios more and more and it will be necessary in the future to go beyond the simplistic one-zone SSC paradigm. In particular, the investigation of additional intrinsic effects in multi-zone scenarios, which would require describing and constraining light travel effects, and in hadronic or lepto-hadronic models (Sol & Zech, 2022) will be a crucial development. The evaluation of Euclidean distance and HID methods in these more complex scenarios will still be needed. However, multi-zone scenarios and hadronic processes might face difficulties in reproducing fast-varying flares. Beyond these possibilities, alternative interesting proposals suggest that some flares could be due to pure geometrical phenomena inducing sudden changes in the Doppler factor of the emitting zone. Such models were developed in the literature for a few blazars showing periodicities or helical jet structures (Mohan & Mangalam, 2015; Raiteri et al., 2017; Sobacchi et al., 2017; Sarkar et al., 2020; Roy et al., 2022; Abe et al., 2024; Chen et al., 2024). This type of flares would certainly deserve specific studies in the context of LIV searches, since they could provide an initial signal with possibly limited intrinsic features, or at least different from standard SSC flares.
@@
Future VHE and X-ray instruments will eventually shed new light on this crucial quest. In particular, the development of multi-wavelength and multi-messenger astronomy and the advent of the Cherenkov Telescope Array (CTA) at VHE will provide for the first time a significant sample of blazar flares observed with good spectral and temporal resolutions and coverage, at different redshifts. This should firmly constrain the true nature of blazar flares, allow discriminating intrinsic energy-dependent time delays from LIV-induced ones, and usher in a new era for both LIV and flare modelling research.